# Self-Supervised Skeleton-Based Action Representation Learning: A Benchmark and Beyond

### v2

@@ 2.2.3 Contrastive-Learning Methods
Contrastive learning has been proven effective for different data modalities, e.g., images, point cloud, as well as skeleton data. Generally, contrastive learning pursues the consistency of the positive samples, which are usually the augmented counterparts of the original data. MoCo [26] utilizes the negative samples to establish an instance discrimination pretext task. Meanwhile, self-distillation [22] and feature decoupling [93, 2] concepts have also been explored in contrastive learning. These early pioneering works [26, 8, 22, 93], have made a huge impact and encouraged unique designs for skeleton contrastive learning.
For skeleton contrastive representation learning, most existing methods are based on MoCo v2 [26]. It adopts an asynchronously momentum-updated key encoder and an online query encoder, along with a memory queue to store a large number of consistent negative samples. Specifically, different augmentations are applied to the skeleton $x$ to generate the positive pair $(x_{q},x_{k})$, while the negative sample features $m_{i}$ are stored in a memory queue. The model is constrained to retrieve the positives among the negative samples, optimizing the following InfoNCE objective $\mathcal{L}_{Info}(z_{q},z_{k})$:
@@
where the $z_{q}$/$z_{k}$ is the query/key embedding encoded by the query/key model and $\tau$ is the temperature hyper-parameter. $m_{i}$ is the $i$-th feature anchor as the negative sample. Based on this, we review current skeleton contrastive learning methods from three aspects of design, i.e., augmentation design and learning strategy, objective engineering, and cross-modal learning.
1) Augmentation Design and Learning Strategy. Data augmentation exposes novel motion patterns and generates diverse positive views, which have been found crucial to the success of contrastive learning [73, 24]. Different from images, data augmentation for skeleton is relatively less developed. Therefore, the earlier works mainly focus on exploring practical spatial-temporal augmentations for skeleton contrastive learning [60, 72, 20]. For example, Rao et al. [60] proposed a series of augmentations including Rotation, Shear, Reverse, Gaussian Noise, Gaussian Blur, Joint Mask and Channel Mask, some of which are used as the default basic augmentations in the future research. Notably, the optimal augmentations for different backbones are usually different. For example, the Joint Mask is found detrimental for GCNs [97], while beneficial for GRU model [72], which implies that GCNs are more sensitive to the spatial corruption. Generally, this difference arises on account of different modeling mechanisms and model capacities, which further increases the difficulty of the augmentation design and selection.
Inspired by the success of mixing-based augmentations for images, Chen et al. [94] proposed SkeleMixCLR equipped with SkeleMix augmentation which combines the topological information of different skeleton sequences. To learn from this mixed skeleton, SkeleMixCLR obtains the corresponding part level and the whole-body level features and pursues the local-global invariance. Based on this, SkeAttnCLR [29] further designs an attention mechanism to perform local contrastive learning on salient and non-salient features. Due to the generated novel input views and regularization effect on the feature space, mixing augmentation often leads to consistent improvement for representations.
To further improve consistency learning, researchers have made endeavors to introduce more and stronger data augmentations. However, the ensuing problem is the over-distortion [1] of the augmented data, leading to the model performance degradation. In other words, some strong augmentations would seriously corrupt the semantic information, change the data distribution, and result in the difficulty of model consistency learning. To address this, AimCLR [24] utilizes two branches to encode the weakly and strongly augmented views, respectively. Then the model optimizes the similarity of the distributions output by two branches as a soft consistency learning target:
$\mathcal{L}_{Soft}$ $=KL(p(z|z_{weak})\|p(z|z_{strong})),$ (3)
@@
where $z_{weak}$ and $z_{strong}$ are the corresponding embeddings of weakly and strongly augmented views. Instead of the one-hot target, this objective utilizes the similarity distribution as the soft target to guide consistency learning, which can be viewed as a self-distillation process. Further, Zhang et al. [97] introduced more strong augmentations, e.g., randomly dropping the skeleton edges and joints, and proposed a hierarchical contrastive learning framework. It performs a decoupled progressive augmentation invariance learning by optimizing the consistency only between the augmented samples with adjacent strength, where the weakly augmented branch serves as the mimic target of the strongly augmented branch. These works show that the weakly augmented views can effectively guide the learning of the corresponding strongly augmented samples, leading to more stable representation learning and improvement.
In addition, Lin et al. [41] explicitly distinguished the static regions and motion regions, namely, actionlet, in human skeleton, and introduced a motion-aware augmentation strategy. By mining the actionlet in an unsupervised manner, the semantic-reserving augmentations are employed for actionlet regions, while the noise perturbations for non-actionlet regions, avoiding the over-distortion problem. On the other hand, instead of augmenting at the input level, HaLP [61] proposes a latent positive hallucinating method by exploring the latent space around the corresponding prototype.
@@
2) Objective Engineering. In addition to the widely used InfoNCE loss, some SSL works have also made efforts to explore new loss functions for extra regularization or new objectives. Here we introduce them from the perspective of False Negative (FN) problem, which widely exists in skeleton contrastive learning based on negative examples, e.g., MoCo v2. Concretely, false negatives refer to the negative samples but from the same semantic category. Traditional contrastive learning relies on the one-hot label and directly pushes them away in the feature space, which forces the model to discard the shared semantic information and leads to slow convergence [30]. We point out that, due to the lack of description of objects and backgrounds, there are fewer action categories represented by skeleton data, which leads to a more serious FN problem. To this end, the following aspects are considered to tackle this issue.
The first straightforward solution is based on the false negative detection and cancellation [38, 24, 96]. These methods first calculate the similarity between the sample and negatives, selecting the top $k$ negatives that are most similar as the potential false negatives, which are then involved as extended positives in contrastive learning.
Some other methods [52, 98] suggest employing the distillation objective as an adaptive re-weighting regularization for the one-hot instance discrimination pretext task, which can be formulated as follows:
$\mathcal{L}_{KD}$ $=-{p}\left({z}_{k},\tau_{k}\right)\log{p}\left(z_{q},\tau_{q}\right),$ (4)
@@
Besides, prototype-based contrastive learning is also studied for high-level semantic consistency learning. It usually performs clustering in feature space to assign instances to different cluster prototypes as a pseudo-semantic label. Then the model learns more high-level semantics by contrasting different prototypes. By virtue of such a way, the model pays more attention to the cluster-level discrimination task rather than the instance-level, which alleviates the false negative problem.
In addition to the above designs, as introduced above, some negative-sample-free contrastive frameworks have been developed recently, which avoid the difficulty of explicitly specifying negative examples. Based on BYOL [22], the works [17, 7] performs the positive-only consistency learning in the hyperbolic space. Some other works [95, 70], inspired by the feature decorrelation concept, propose to learn the decorrelated representations based on Barlow twins [93] and Variance-Invariance-Covariance Regularization (VICReg) [2] as the framework.
@@
3) Cross-Modal Learning. Skeleton naturally provides different data modalities to represent human movement, e.g., motion and bone modality defined in the previous works [63, 10]. Meanwhile, many works [63, 10] have proved that the fusion of different modal knowledge can be beneficial to achieve richer and more representative action depictions. Therefore, cross-modal learning has attracted the attention of researchers for skeleton representation learning.
@@
CRRL [78] performs contrastive learning between the joint and motion modalities, where the encoders for motion and joint data are homologous, i.e., obtained by momentum-updated strategy. However, this design can be unreasonable because the model can have difficulty dealing with two modalities simultaneously. In contrast, CrosSCLR [38] adopts three separate encoders for joint, bone, and motion data, and aims to align the distributions among the neighborhoods in latent space. Intuitively, it encourages the sample embedding together with its neighbors of modality $v$ should also be close in the latent space of modality $u$. However, this leads to a two-stage training process, as the model needs to be pre-trained on a single view first to obtain reliable unique modal knowledge. To this end, CMD [52] proposes a general bidirectional distillation objective, where the modal knowledge is directly modeled as the whole similarity distribution in the customized latent space. The proposed cross-modal distillation objective and the instance discrimination task jointly optimize the model, yielding a concise single training phase. On the other hand, to get rid of the limitation of using separate encoders for different modalities, UmURL [70] develops a unified modality-agnostic encoder, which can handle different modal inputs relying on the modality-specific input embedding layer and feature projection layer.
@@
Meanwhile, we can also organize skeleton data into different representations as shown in Fig. 4, e.g., the temporal series, 2D pseudo-images with frame and joint dimensions, and spatial-temporal graphs with nodes and edges. Following this concept, ISC [72] encodes these different skeleton representations with different model backbones, which are then projected into a shared latent space. A cross-modal contrastive learning loss is applied for pre-training.
@@
4) Others. In addition to the aforementioned literature, some works also employ contrastive learning. Considering the hierarchical structure of skeleton data, HiCo [16] performs contrastive learning in a hierarchical manner to model features of different levels. Some works [40, 83, 71, 78] combine the contrastive learning paradigm with others in a multi-tasking manner. Notably, in addition to simply combining reconstruction and contrastive learning pretext tasks, PCM3 [98] proposes a collaborative design to further improve the representation learning. Eq-Contrast [42] formulates the temporal sorting task into an equivariant contrastive learning objective for multi-task pre-training. Meanwhile, not limited to SSL, contrastive learning is also studied in supervised [101] and semi-supervised [81, 65] learning for skeleton.
@@ 2.3 Summary and Discussion
The skeleton-based SSL literature is categorized into context-based, generative learning, and contrastive learning types. All these methods aims to capture joint features as well as their relationships from unlabeled skeleton data, to obtain a meaningful representation of the motion. Due to the lack of ground-truth labels, these methodologies introduce prior knowledge in different means to enable representation learning.
Specifically, context-based methods introduce the prior knowledge by the predefined pretext-task designs, e.g., predicting the equivariant properties of some specific transformation, learning the spatial concurrence or temporal dependencies of skeleton joints. Generative approaches generally reconstruct or predict the skeleton target without much prior information. However, some recent works have shown that some special designs, e.g., masking the motion regions, can further boost the performance, where the prior knowledge is also utilized that the moving parts can better represent the action. In the context of contrastive learning, the prior information is mostly reflected through the data augmentations for the construction of positive pairs, which affect the invariance learning of transformation-related features. Overall, although downstream tasks are often unknowable in pre-training, these methods can obtain a potentially beneficial representation space by introducing different prior knowledge.
@@ 3.1 Motivation
By reviewing the previous works, we find that most existing skeleton-based SSL methods employ a single paradigm or simply combine different methods, evaluated on the action recognition task solely. We argue this ignores the hierarchical design for representation learning of different granularity and leaves the generalization capacity of model leashed, as we will verify later in Sec. 4.7. To construct a strong baseline model for skeleton-based SSL, we propose prompted contrast with masked motion modeling (PCM3++), which combines contrastive learning and MSM by exploring the potential collaboration. It enables versatile representation learning by joint, clip, and sequence level feature modeling, and significantly improves the performance of different downstream tasks.
Specifically, our method is based on our previous work [98] and integrates valuable improvement for better versatile representation learning. We first briefly introduce the joint-level and sequence-level feature modeling of the baseline in Sec. 3.2 - 3.4. Based on this, we introduce a novel asymmetric clip-level contrastive learning method in Sec. 3.2. The training strategy is presented in Sec. 3.5, including the newly proposed post-distillation policy for further representation refinement. The readers can refer to [98] for more details.

### v3

@@ 2.2.3 Contrastive-Learning Methods
Contrastive learning has been proven effective for different data modalities, e.g., images, point cloud, as well as skeleton data. Generally, contrastive learning pursues the consistency of the positive samples, which are usually the augmented counterparts of the original data. MoCo [he2020momentum] utilizes the negative samples to establish an instance discrimination pretext task. Meanwhile, self-distillation [grill2020bootstrap] and feature decoupling [zbontar2021barlow, bardes2021vicreg] concepts have also been explored in contrastive learning. These early pioneering works [he2020momentum, chen2020simple, grill2020bootstrap, zbontar2021barlow, robinson2020contrastive], have made a huge impact and encouraged unique designs for skeleton contrastive learning.
For skeleton contrastive representation learning, most existing methods are based on MoCo v2 [he2020momentum]. It adopts an asynchronously momentum-updated key encoder and an online query encoder, along with a memory queue to store a large number of consistent negative samples. Specifically, different augmentations are applied to the skeleton $x$ to generate the positive pair $(x_{q},x_{k})$, while the negative sample features $m_{i}$ are stored in a memory queue. The model is constrained to retrieve the positives among the negative samples, optimizing the following InfoNCE objective $\mathcal{L}_{Info}(z_{q},z_{k})$:
@@
where the $z_{q}$$z_{k}$ is the query/key embedding encoded by the query/key model and $\tau$ is the temperature hyper-parameter. $m_{i}$ is the $i$-th feature anchor as the negative sample. Based on this, we review current skeleton contrastive learning methods from three aspects of design, i.e., augmentation design and learning strategy, objective engineering, and cross-modal learning.
1) Augmentation Design and Learning Strategy. Data augmentation exposes novel motion patterns and generates diverse positive views, which have been found crucial to the success of contrastive learning [tian2020makes, guo2022aimclr]. Unlike images, data augmentation for skeleton is relatively less developed. Therefore, the earlier works mainly focus on exploring practical spatial-temporal augmentations for skeleton contrastive learning [rao2021augmented, thoker2021skeleton, gao2021efficient]. For example, Rao et al. [rao2021augmented] proposed a series of augmentations including Rotation, Shear, Reverse, Gaussian Noise, Gaussian Blur, Joint Mask and Channel Mask, some of which are used as the default basic augmentations in the future research. Notably, the optimal augmentations for different backbones are usually different. For example, the Joint Mask is found detrimental for GCNs [zhang2022hiclr], while beneficial for GRU model [thoker2021skeleton], which implies that GCNs are more sensitive to the spatial corruption. Generally, this difference arises on account of different modeling mechanisms and model capacities, which further increases the difficulty of the augmentation design and selection.
Inspired by the success of mixing-based augmentations for images, Chen et al. [chen2022skelemixclr] proposed SkeleMixCLR equipped with SkeleMix augmentation which combines the topological information of different skeleton sequences. To learn from this mixed skeleton, SkeleMixCLR obtains the corresponding part level and the whole-body level features and pursues the local-global invariance. Based on this, SkeAttnCLR [hua2023part] further designs an attention mechanism to perform local contrastive learning on salient and non-salient features. Due to the generated novel input views and regularization effect on the feature space, mixing augmentation often leads to consistent improvement for representations.
To further improve consistency learning, researchers have made endeavors to introduce more and stronger data augmentations. However, the ensuing problem is the over-distortion [bai2022directional] of the augmented data, leading to the model performance degradation. In other words, some strong augmentations would seriously corrupt the semantic information, change the data distribution, and result in the difficulty of model consistency learning. To address this, AimCLR [guo2022aimclr] utilizes two branches to encode the weakly and strongly augmented views, respectively. Then the model optimizes the similarity of the distributions output by two branches as a soft consistency learning target:
$\mathcal{L}_{Soft}(z)$ $=KL(p(z|z_{weak})\|p(z|z_{strong})),$ (3)
@@
where $z_{weak}$ and $z_{strong}$ are the corresponding embeddings of weakly and strongly augmented views. Instead of the one-hot target, this objective utilizes the similarity distribution as the soft target to guide consistency learning, which can be viewed as a self-distillation process. Further, Zhang et al. [zhang2022hiclr] introduced more strong augmentations, e.g., randomly dropping the skeleton edges and joints, and proposed a hierarchical contrastive learning framework. It performs a decoupled progressive augmentation invariance learning by optimizing the consistency only between the augmented samples with adjacent strength, where the weakly augmented branch serves as the mimic target of the strongly augmented branch. These works show that the weakly augmented views can effectively guide the learning of the corresponding strongly augmented samples, leading to more stable representation learning and improvement.
In addition, Lin et al. [lin2023actionlet] explicitly distinguished the static regions and motion regions, namely, actionlet, in human skeleton, and introduced a motion-aware augmentation strategy. By mining the actionlet in an unsupervised manner, the semantic-reserving augmentations are employed for actionlet regions, while the noise perturbations for non-actionlet regions, avoiding the over-distortion problem. On the other hand, instead of augmenting at the input level, HaLP [shah2023halp] proposes a latent positive hallucinating method by exploring the latent space around the corresponding prototype.
@@
2) Objective Engineering. In addition to the widely used InfoNCE loss, some SSL works have also made efforts to explore new loss functions for extra regularization or new objectives. Here we introduce them from the perspective of False Negative (FN) problem, which widely exists in skeleton contrastive learning based on negative examples, e.g., MoCo v2. Concretely, false negatives refer to the negative samples but from the same semantic category. Traditional contrastive learning relies on the one-hot label and directly pushes them away in the feature space, which forces the model to discard the shared semantic information and leads to slow convergence [huynh2022boosting]. We point out that, due to the lack of description of objects and backgrounds, there are fewer action categories represented by skeleton data, which leads to a more serious FN problem. To this end, the following aspects are considered to tackle this issue.
The first straightforward solution is based on the false negative detection and cancellation [chen2021incremental, huynh2022boosting, li20213d, guo2022aimclr, zhang2022cpm]. These methods first calculate the similarity between the sample and negatives, selecting the top $k$ negatives that are most similar as the potential false negatives, which are then involved as extended positives in contrastive learning.
Some other methods [xie2022delving, mao2022cmd, zhang2023prompted] suggest employing the distillation objective as an adaptive re-weighting regularization for the one-hot instance discrimination pretext task, which can be formulated as follows:
$\mathcal{L}_{KD}(z_{q},z_{k})$ $=-{p}\left({z}_{k},\tau_{k}\right)\log{p}\left(z_{q},\tau_{q}\right),$ (4)
@@
Besides, prototype-based contrastive learning is also studied for high-level semantic consistency learning [li2020prototypical, caron2020unsupervised, xu2021prototypical]. It usually performs clustering in feature space to assign instances to different cluster prototypes as a pseudo-semantic label. Then the model learns more high-level semantics by contrasting different prototypes. By virtue of such a way, the model pays more attention to the cluster-level discrimination task rather than instance-level to alleviate the false negative problem.
In addition to the above designs, as introduced above, some negative-sample-free contrastive frameworks have been developed recently, which avoid the difficulty of explicitly specifying negative examples. Based on BYOL [grill2020bootstrap], the works [francohyperbolic, chen2023self] performs the positive-only consistency learning in the hyperbolic space. Some other works [zhang2022skeletal, sun2023unified], inspired by the feature decorrelation concept, propose to learn the decorrelated representations based on Barlow twins [zbontar2021barlow] and Variance-Invariance-Covariance Regularization (VICReg) [bardes2021vicreg] as the framework.
@@
Figure 5: Summary of the advantages/limitations of different skeleton-based SSL methodologies.
@@
3) Cross-Modal Learning. Skeleton naturally provides different data modalities to represent human movement, e.g., motion and bone modality defined in the previous works [shi2019two, chen2021channel]. Meanwhile, many works [shi2019two, chen2021channel] have proved that the fusion of different modal knowledge can be beneficial to achieve richer and more representative action depictions. For example, joint contains essential information about human movements while lacking the explicit encoding of some high-order information. Bone view can capture the structural patterns well. However, it is subject to the predefined human model and still lacks temporal information. Motion explicitly encodes the temporal dynamics while losing the spatial information. It is promising to explore these different modalities to achieve better action representation.
@@
CRRL [wang2022contrast] performs contrastive learning between the joint and motion modalities, where the encoders for motion and joint data are homologous, i.e., obtained by momentum-updated strategy. However, this design can be unreasonable because the model can have difficulty dealing with two modalities simultaneously. In contrast, CrosSCLR [li20213d] adopts three separate encoders for joint, bone, and motion data, and aims to align the distributions among the neighborhoods in latent space. Intuitively, it encourages the sample embedding together with its neighbors of modality $v$ should also be close in the latent space of modality $u$. However, this leads to a two-stage training process, as the model needs to be pre-trained on a single view first to obtain reliable unique modal knowledge. To this end, CMD [mao2022cmd] proposes a general bidirectional distillation objective, where the modal knowledge is directly modeled as the whole similarity distribution in the customized latent space. The proposed cross-modal distillation objective and the instance discrimination task jointly optimize the model, yielding a concise single training phase. On the other hand, to get rid of the limitation of using separate encoders for different modalities, UmURL [sun2023unified] develops a unified modality-agnostic encoder, which can handle different modal inputs relying on the modality-specific input embedding layer and feature projection layer.
@@
Meanwhile, we can also organize skeleton data into different representations as shown in Fig. 4, e.g., the temporal series, 2D pseudo-images with frame and joint dimensions, and spatial-temporal graphs with nodes and edges. Following this concept, ISC [thoker2021skeleton] encodes these different skeleton representations with different model backbones, which are then projected into a shared latent space. A cross-modal contrastive learning loss is applied for pre-training.
@@
4) Others. In addition to the aforementioned literature, some works also employ contrastive learning. Considering the hierarchical structure of skeleton data, HiCo [hico2023] performs contrastive learning in a hierarchical manner to model features of different levels. Some works [lin2020ms2l, xu2021prototypical, tanfous2022and, wang2022contrast] combine the contrastive learning paradigm with others in a multi-tasking manner. Notably, in addition to simply combining reconstruction and contrastive learning pretext tasks, PCM3 [zhang2023prompted] proposes a collaborative design to further improve the representation learning. Eq-Contrast [lin2024mutual] formulates the temporal sorting task into an equivariant contrastive learning objective for multi-task pre-training. Meanwhile, not limited to SSL, contrastive learning is also studied in supervised [zhou2023learning] and semi-supervised [xu2022x, shu2022multi] learning for skeleton.
@@ 2.3 Summary and Discussion
The skeleton-based SSL literature is categorized into context-based, generative learning, and contrastive learning types. All these methods aims to capture joint features as well as their relationships from unlabeled skeleton data, to obtain a meaningful representation of the motion. Due to the lack of ground-truth labels, these methodologies introduce prior knowledge in different means to enable representation learning, e.g., context task design, augmentation choices, and masking strategy. To summary the different skeleton-based SSL methods based on our taxonomy, we present the advantages and limitations of each methodology category in Fig. 5.
Context-based methods are often simple to implement, flexible for design, and training-friendly compared with the other two paradigms. However, as evidenced by the previous literature, the main concern is the limited model capacity of generalization and semantic understanding due to the gap between the low-level pretext tasks and the downstream tasks. Contrastive learning is recognized by its highly discriminative representations with strong performance, especially under linear evaluation protocol. More importantly, it has been proven effective for different backbones for skeleton modeling. The other side of the coin is augmentation sensitivity, false negative problem, and high training cost as previously discussed. Meanwhile, some researchers point out the concern of dimensional collapse in contrastive learning [jingunderstanding, lin2024mutual], where the embedding vectors only span a lower dimensional space, limiting the representation capacity. Finally, the masked skeleton modeling is discussed considering its dominant position in generative learning. It is well compatible with Transformer, with expected strong generalization performance, especially under fine-tuning protocol. However, performance can significantly drop when employing other backbones, which can limit its deployment in some low-latency, resource-constrained scenarios. Meanwhile, it is also concerned about masking strategy sensitivity and time-consuming training.
To end this section and possibly motivate future work, we further summarize and emphasize the special design of skeleton-based SSL approaches that go beyond the high-level SSL methodology taxonomy.
$\bullet$ Context Pretext Tasks Based on Human Topology and Kinematics. Different from images, skeleton pretext tasks focus on the spatial-temporal data structure learning of human body, e.g., temporal sorting, angle, and displacement prediction.
$\bullet$ Data Augmentation for Novel Motion Patterns. Skeleton augmentation has always been the focus to boost contrastive learning. These efforts include the spatial and temporal, invariant and equivariant [lin2024mutual], weak and strong [guo2022aimclr, zhang2022hiclr], intra and inter augmentations [chen2022skelemixclr, zhang2023prompted] as well as the corresponding learning strategy.
$\bullet$ Cross-Modal Knowledge Modeling. Skeleton naturally provides different data modalities (or views) to represent human movement, e.g., joint, motion, and bone, enabling cross-modal knowledge modeling for better action representations.
$\bullet$ False-Negative Cancellation. As discussed, the diversity of human motions is much less than natural images, leading to a more serious false negative problem, which is an important topic in skeleton contrastive learning.
$\bullet$ Masking Strategy in MSM. As a crucial component in MSM, spatial-temporal masking strategy has also been widely studied specifically for the highly structured human skeleton data.
$\bullet$ Multi-Granularity Representation Learning. Skeleton data naturally presents a hierarchy, e.g., sequence-level, clip-level, and joint-level. It can be beneficial to capture these action clues at different scales, enriching the action representation and alleviating the over-fitting problem.
@@ 3.1 Motivation
By reviewing the previous works, we find that most existing skeleton SSL methods employ a single paradigm or simply combine different methods with action recognition task for evaluation solely, which limits the model capacity of learning representations at different granularity and levels and leaves the generalization capacity leashed. To this end, we propose prompted contrast with masked motion modeling (PCM3++), which combines contrastive learning and MSM by exploring potential collaboration. It enables versatile representation learning by joint, clip, and sequence level feature modeling, and significantly improves the performance of different downstream tasks.
Table 1: Main notations and the definition.
Notation Definition
$s^{\prime}(s_{ach})$, $z^{\prime}(z_{ach})$ The global key-branch intra-augmented
skeleton and its embedding.
$s_{intra}$, $z_{intra}$ The query-branch intra-augmented
skeleton and its embedding.
$s_{inter}$, $z_{inter}$ The query-branch inter-augmented
skeleton and its embedding.
$s_{clip}$, $z_{clip}$ The query-branch temporal-augmented
skeleton clip and its embedding.
$s_{mask}$, $z_{mask}$ The query-branch masked
skeleton and its embedding.
$s_{predict}$, $z_{predict}$ The reconstructed skeleton from $s_{mask}$
and its embedding.
$z_{inter}^{\prime}$ The linear interpolation of two skeleton
embeddings in mix inter-augmentation.
$p_{*}$ The prompts of different views.
$f(\cdot),h(\cdot),dec(\cdot)$ The encoder, projector, and decoder.
Specifically, our method is based on our previous work [zhang2023prompted], which combines contrastive learning and masked skeleton modeling in a synergetic manner. Based on this, we further extend and strengthen the design with the key idea, i.e., versatile action representation learning. In the following, we first introduce the contrastive learning (Sec. 3.2) and MSM (Sec. 3.3) scheme involved in baseline method as well as their collaboration design (Sec. 3.4). Based on this, we introduce our new improvements in PCM3++ compared to [zhang2023prompted]. Concretely, an effective clip-level contrastive learning method is designed in Sec. 3.2.(3), which significantly boosts the capacity of short-term motion modeling as illustrated in Fig. 6. Besides, a post-distillation strategy is developed in Sec. 3.5.(2) to further boost the representation compactness. These advancements further improve the representation generalization for various downstream tasks. For clarity, the main notations are summarized in Table 1.